[Brahms-l] Shift report 20040224 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Tue Feb 24 2004 - 15:59:34 EST
------------------------------------------------------------
Date: Tue Feb 24 15:59:34 2004, Shift: 08:00-16:00
Supervisor: BHS, with: 
Log book#15 pages 66 - 66

Shift Summary:
------------------------------------------------------------
1: Machine had one long store 08.00-12.30, then lost beam and  has been down with power supply problems since then. An access is still under ways.

2: Data runs with FS@8, 1/2A, MRS@40, 6A: 10578-10581. 197k trig2, 700k trig3.

3: No problems at all. A lady came by inspecting the counting house floors and said that they would be waxed and cleaned :-)

------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Feb 24 16:00:14 2004

This archive was generated by hypermail 2.1.8 : Tue Feb 24 2004 - 16:00:35 EST